# The Role of Entanglement in Quantum Reservoir Computing with Coupled Kerr Nonlinear Oscillators

### v1

@@ Abstract
Quantum Reservoir Computing (QRC) uses quantum dynamics to efficiently process temporal data. In this work, we investigate a QRC framework based on two coupled Kerr nonlinear oscillators, a system well-suited for time-series prediction tasks due to its complex nonlinear interactions and potentially high-dimensional state space. We explore how its performance in time-series prediction depends on key physical parameters: input drive strength, Kerr nonlinearity, and oscillator coupling, and analyze the role of entanglement in improving the reservoir’s computational performance, focusing on its effect on predicting non-trivial time series. Using logarithmic negativity to quantify entanglement and normalized root mean square error (NRMSE) to evaluate predictive accuracy, our results suggest that entanglement provides a computational advantage on average—up to a threshold in the input frequency—that persists under some levels of dissipation and dephasing. In particular, we find that higher dissipation rates can enhance performance. While the entanglement advantage manifests as improvements in both average and worst-case performance, it does not lead to improvements in the best-case error. These findings contribute to the broader understanding of quantum reservoirs for high performance, efficient quantum machine learning and time-series forecasting.
@@ I Introduction
A key question driving QRC research is what intrinsic quantum features can provide computational advantages over classical reservoirs. Quantum systems offer two notable features: (i) a state space that grows exponentially with the number of constituents, and (ii) uniquely quantum phenomena such as superposition and entanglement. The former means that a moderately sized quantum reservoir can explore a vast space of configurations, which in principle could embed very complex mappings of the input data [26]. The latter – entanglement and related non-classical correlations – might endow the reservoir with richer dynamics and memory than any classical counterpart [26].
While some studies show that quantum reservoir computing can function without strongly non-classical states[27], others suggest that genuine quantum resources—such as entanglement, coherence, or quantum chaos—may offer enhanced computational power. For example, memory capacity in qubit-based reservoirs has been found to correlate with entanglement, and recent work showed that this advantage depends on the input frequency [26, 28]. However, more “quantumness” is not always better: studies have shown that in certain tasks, such as NARMA, introducing squeezing (a form of nonclassicality) can reduce performance [29]. These findings underscore that while quantum effects can be beneficial, their impact on reservoir computing remains subtle and task-dependent—especially the role of entanglement, which is still not fully understood.
This paper aims to expand this area of investigation. In contrast to the previous studies mentioned above, our work explicitly focuses on entanglement in a reservoir formed by two coupled Kerr oscillators and its relationship with predictive performance. By varying the system parameters such as coupling, input (driving) strength, and nonlinearity, we can tune the reservoir’s entanglement, which we quantify by means of logarithmic negativity, while evaluating the prediction performance on the target time series by calculating normalized root mean square error (NRMSE). We also explore how the relationship is affected by dissipation and dephasing. Notably, as we focus on low-excitation-number dynamics, our analysis is performed in the fully quantum regime.
@@
Figure 1: Overview of the coupled Kerr Oscillator QRC. The oscillators are driven by the input signal with amplitudes $\epsilon_{a}$ and $\epsilon_{b}$ at a coupling strength of $g$. The reservoir state is measured in the Fock basis (read-out) at moments of $t=\delta t/m$. In response, the quantum state of the system evolves, and the computational nodes evolve due to the natural dynamics of the system and input signal. The state of the system is measured in the Fock basis at different times at a constant rate (sampling via multiplexing). The output of the reservoir is fed to the ridge regression to generate weights. The weights are then used to generate the output time series, which is a prediction of the target.
@@ II.1 Physical Model
We model the coupled Kerr oscillator system as an open quantum system with Markovian dynamics. The evolution is governed by a Lindblad master equation for the density operator $\rho(t)$:
$\frac{d\hat{\rho}(t)}{dt}=i[\hat{H}(t),\hat{\rho}(t)]+\sum_{i=a,b}(\hat{c}_{i}\hat{\rho}(t)\hat{c}_{i}^{\dagger}-\frac{1}{2}\{\hat{c}_{i}^{\dagger}\hat{c}_{i},\hat{\rho}(t)\})$ (5)
where the collapse operators describing decay in modes $a$ and $b$ are given by $c_{a}=\sqrt{2\kappa_{a}}\hat{a}+\sqrt{\kappa_{\phi}}\hat{N}_{a}$ and $c_{b}=\sqrt{2\kappa_{b}}\hat{b}+\sqrt{\kappa_{\phi}}\hat{N}_{b}$ with $\kappa_{\phi}$ being associated dephasing rate. Throughout this work, we set the parameters to be equal for both oscillators and, for convenience, we work in a rotating frame at the drive frequencies (taken equal to $\omega_{a}$ and $\omega_{b}$ for each mode); in this frame the free oscillation terms are eliminated (effectively setting $\omega_{a}=\omega_{b}=0$). It is worth noting that $\hat{H}_{int}$ does not include nonlinear coupling proportional to $\hat{N}_{a}\hat{N}_{b}$, but only the linear one, described by Eq. 3. The same Hamiltonian $\hat{H}_{nl}+\hat{H}_{int}$ was used, e.g., by Bernstein [33] and Chefles and Barnett [34] to describe the nonlinear coupler. Note that we are using dimensionless units throughout this study.
@@ II.2 Learning and Simulation
The first half of the input time series is used for training and the other half for testing. For training, Tikhonov regularization is utilized with a regularization parameter that is chosen to get maximum performance. For quantifying the performance of the reservoir, we use the normalized root mean square error (NRMSE), which is calculated as
@@ III Results
Figure 2: Time evolution of entanglement (top row) and corresponding time-series prediction performance (bottom row) at different input frequencies $f$. Top: Behavior of logarithmic negativity in time during testing and training. To ensure the reservoir reflects stable dynamics rather than initial transients, we use the portion of the evolution after the vertical dashed line for training and testing. Bottom: Examples of input sequences at three different frequency scales and The corresponding predictions of the reservoir at coupling rate of $g=0.9$ , input strengths of $\epsilon_{a}=\epsilon_{b}=3$, nonlinearities of $K_{a}=K_{b}=1$ and dissipation rates of $\kappa_{a}=\kappa_{b}=0.1$ in a rotating frame. The performance gets worse as the input frequency increases.
@@ III.1 Input Strength
We randomly selected 30 input strengths ($\epsilon$) from the interval $[0,10]$ and evaluated the normalized root mean square error (NRMSE) for input time series at various input frequency scales while computing the logarithmic negativity of the system in each instance. It is important to note that we utilize the same 30 values of input strengths for all frequencies. Furthermore, we ensure that the average excitation number remains below one as discussed in section II.4 and appendix B. The outcome is presented in Fig. 3.
@@ III.2 Kerr Nonlinearity
We randomly selected 30 Kerr nonlinearity values ($K$) from the interval $[0,10]$ and evaluated the NRMSE for input time series at different input frequency scales while calculating the logarithmic negativity of the system in each case. Note that we use the same 30 values of $K$ for all frequencies. The results are shown in Fig. 4.
@@ III.3 Coupling Constant
We randomly selected 30 coupling constant values ($g$) from the interval $[0,10]$ and evaluated the NRMSE on input time series at different input frequency scales while calculating the logarithmic negativity of the system in each case. Note that we use the same 30 values of $g$ for all frequencies. The results are shown in Fig. 5.
@@ III.4 Entanglement Advantage
Figure 6: Prediction error (NRMSE) vs. entanglement in the case of zero dephasing ($\kappa_{\phi}=0$) and a dissipation rate of $\kappa_{a}=\kappa_{b}=0.1$, for different input frequencies. The data from figs. 3, 4 and 5 are combined and binned based on entanglement levels, quantified via logarithmic negativity, with vertical error bars representing the standard deviation within each bin. Left: Minimum NRMSE within each entanglement bin. No consistent correlation is observed. Middle: Mean NRMSE across all configurations in each bin. This captures the overall performance trend with respect to entanglement. Moderate entanglement improves performance on average. There is an optimum level of entanglement. However, a threshold in the frequency is observed at $f=2.5$ where no trend exists. Right: Maximum NRMSE per entanglement bin. This reflects the worst-performing cases within each bin. A stronger relation is seen here: higher entanglement reduces the worst-case error more than the average error, but again, a frequency threshold is observed.
@@ III.5 Role of Dissipation
To examine the impact of dissipation, we repeat the data generation process and binned analysis described earlier in the presence of different dissipation rates, $\kappa_{a}=\kappa_{b}=\kappa$ and compute the average entanglement and NRMSE to produce the same type of plot as the middle panel of Fig. 6, now for different $\kappa$ (see Fig. 9 in Appendix A). It is worth mentioning that our system does not reach a steady state in the absence of dissipation and therefore does not exhibit steady-state entanglement; hence, we do not consider the zero-dissipation case. To better analyze the effect of dissipation, we extract the optimal points (those with minimum error) and plot their corresponding NRMSE and entanglement as a function of dissipation for different input frequencies, resulting in Fig. 7.
As shown in Fig. 7, a higher dissipation level is better for frequencies below the frequency threshold up to a point where it saturates. By looking at Fig. 9 and Appendix A one can see that realizing an entanglement advantage requires sufficient dissipation, with slightly higher values offering a consistent advantage across all frequency scales below the frequency threshold.
@@ III.6 Role of Dephasing
To examine the impact of dephasing, we repeat the data generation process and binned analysis described earlier in the presence of different dephasing rates $\kappa_{\phi}$, and compute the average entanglement and NRMSE to produce the same type of plot as in Fig. 6, now for different $\kappa_{\phi}$ (Fig. 10 in Appendix A). To more clearly assess the role of dephasing, we identify the optimal points—those corresponding to the minimum prediction error—and plot their associated NRMSE and entanglement as a function of dephasing for different input frequencies, as shown in Fig.8 and Fig. 10 in Appendix A.
@@ IV Discussion and Conclusion
Our findings indicate that within the various system configurations analyzed, entanglement is indeed linked to improvement in both average and worst-case predictive performance, albeit only within a certain range of operational parameters and up to a threshold in the input frequency. It was noted that optimal average performance, measured in terms of NRMSE, generally takes place when the reservoir maintains a moderate level of entanglement rather than the highest, with the worst-case performance following a similar trend. This effect was most evident for lower input frequency scales, where the temporal structure of the input could be effectively retained and processed by the reservoir dynamics. This kind of relation between quantumness and learning capabilities is consistent with other recent findings in the literature [37, 26, 28], which suggest that excessive “quantumness” may saturate the reservoir or introduce dynamical instability, ultimately hampering learning.
Notably, our results demonstrate that the observed connection between entanglement and performance persists under realistic conditions with dissipation and dephasing (look at Appendix A). We showed that increasing dissipation can actually improve performance, up to a point that it saturates. Likewise, we found that the absence of dephasing is generally optimal for performance and entanglement (look at Appendix A); a small amount of dephasing can be tolerated (and at very low input frequency, it even provided a slight benefit).
There are several limitations to note. First, our study was limited to a two-mode Kerr system with a truncated Hilbert space—a limitation that we addressed by imposing a limit on average excitation number (see Section II.4, Appendix B). While this setup is well-suited for exploring foundational aspects of QRC, it restricts both the memory depth and the expressive capacity of the reservoir. Second, we focused on a specific class of input signals and employed a standard linear readout with ridge regression. Although this choice facilitates controlled benchmarking, our conclusions may not directly generalize to more diverse or noisy targets or to different output architectures. Additionally, our quantification of entanglement relied on logarithmic negativity, which does not capture all forms of non-classical correlation, such as Wigner negativity, quantum discord, magic/nonstabilizerness, or coherence.

### v2

@@ Abstract
Quantum Reservoir Computing (QRC) uses quantum dynamics to efficiently process temporal data. In this work, we investigate a QRC framework based on two coupled Kerr nonlinear oscillators, a system well-suited for time-series prediction tasks due to its complex nonlinear interactions and potentially high-dimensional state space. We explore how its performance in forecasting both linear and nonlinear time-series depends on key physical parameters: input drive strength, Kerr nonlinearity, and oscillator coupling, and analyze the role of entanglement in improving the reservoir’s computational performance, focusing on its effect on predicting non-trivial time series. Using logarithmic negativity to quantify entanglement and normalized root mean square error (NRMSE) to evaluate predictive accuracy, our results suggest that entanglement provides a computational advantage on average—up to a threshold in the input frequency—that persists under some levels of dissipation and dephasing. In particular, we find that higher dissipation rates can enhance performance. While the entanglement advantage manifests as improvements in both average and worst-case performance, it does not lead to improvements in the best-case error. These findings contribute to the broader understanding of quantum reservoirs for high performance, efficient quantum machine learning and time-series forecasting.
@@ I Introduction
A key question driving QRC research is what intrinsic quantum features can provide computational advantages over classical reservoirs. Quantum systems offer two notable features: (i) a state space that grows exponentially with the number of constituents, and (ii) uniquely quantum phenomena such as superposition and entanglement. The former means that a moderately sized quantum reservoir can explore a vast space of configurations, which in principle could embed very complex mappings of the input data [14]. The latter – entanglement and related non-classical correlations –may signal the presence of richer dynamics and memory than any classical counterpart [14].
While some studies show that quantum reservoir computing can function without strongly non-classical states [31], others suggest that genuine quantum resources—such as entanglement, coherence, or quantum chaos—may offer enhanced computational power. For example, memory capacity in qubit-based reservoirs has been found to correlate with entanglement, and recent work showed that this advantage becomes more prominent at higher input frequencies [14, 22]. However, it remains unclear if these findings hold for continuous-variable systems or if they are specific to spin-networks and state-injection encoding mechanisms. Notably, evidence suggests that more “quantumness” is not necessarily better. For instance, studies have shown that in certain tasks, such as NARMA, introducing squeezing (a form of nonclassicality) can reduce performance [12]. These findings underscore that while quantum effects can be beneficial, their impact on reservoir computing remains subtle and task-dependent—especially the role of entanglement, which is still not fully understood.
To address these open questions, we investigate a QRC framework based on two coupled Kerr nonlinear oscillators. We employ Hamiltonian encoding, consistent with previous oscillator-based QRC implementations [15, 9, 19], rather than the state injection techniques often utilized in spin-network memory capacity studies [22]. This strategy is particularly advantageous for forecasting as it circumvents the experimental overhead of repeated state preparation and enables highly nonlinear feature mapping. Furthermore, it avoids the control overheads associated with resetting the reservoir’s memory [27], ensuring that entanglement is not totally destroyed each time. Moreover, we do time-series forecasting both on linear multi-frequency signals and nonlinear NARMA time-series, which is substantially different than memory tasks because forecasting necessitates learning the signal’s underlying dynamics rather than merely storing past inputs. Crucially, this approach uncovers a behavior distinct from spin-network QRC [22]: instead of a monotonic performance gain at high frequencies, we observe a frequency threshold, below which entanglement provides the computational advantage.
We further expand the analysis of noise tolerance by explicitly modeling dephasing alongside dissipation. Through a sweep of physical parameters—including drive strength, nonlinearity, and coupling—we analyze the interplay between entanglement and prediction error in the fully quantum regime.
@@
Figure 1: Overview of the coupled Kerr Oscillator QRC. The oscillators are driven by the input signal with amplitudes $\epsilon_{a}$ and $\epsilon_{b}$ at a coupling strength of $g$. The reservoir state is measured in the Fock basis (read-out) at moments of $t=\delta t/m$ at a constant rate (sampling via multiplexing). In response, the quantum state of the system evolves, and the computational nodes evolve due to the natural dynamics of the system and input signal. The output of the reservoir is fed to the ridge regression to generate weights. The weights are then used to generate the output time series, which is a prediction of the target.
@@ II.1 Physical Model
We model the coupled Kerr oscillator system as an open quantum system with Markovian dynamics. The evolution is governed by a Lindblad master equation for the density operator $\hat{\rho}(t)$:
$\frac{d\hat{\rho}(t)}{dt}=-i[\hat{H}(t),\hat{\rho}(t)]+\sum_{i=1}^{4}(\hat{c}_{i}\hat{\rho}(t)\hat{c}_{i}^{\dagger}-\frac{1}{2}\{\hat{c}_{i}^{\dagger}\hat{c}_{i},\hat{\rho}(t)\})$ (5)
where the collapse operators describing decay in modes $a$ and $b$ are given by $\hat{c_{1}}=\sqrt{2\kappa_{a}}\hat{a}$, $\hat{c_{2}}=\sqrt{2\kappa_{b}}\hat{b}$, $\hat{c_{3}}=\sqrt{\kappa_{\phi}}\hat{N}_{a}$, and $\hat{c_{4}}=\sqrt{\kappa_{\phi}}\hat{N}_{b}$ with $\kappa_{\phi}$ being associated dephasing rate. Throughout this work, we set the parameters to be equal for both oscillators and, for convenience, we work in a rotating frame at the drive frequencies (taken equal to $\omega_{a}$ and $\omega_{b}$ for each mode); in this frame the free oscillation terms are eliminated (effectively setting $\omega_{a}=\omega_{b}=0$). It is worth noting that $\hat{H}_{int}$ does not include nonlinear coupling proportional to $\hat{N}_{a}\hat{N}_{b}$, but only the linear one, described by Eq. 3. The same Hamiltonian $\hat{H}_{nl}+\hat{H}_{int}$ was used, e.g., by Bernstein [3] and Chefles and Barnett [7] to describe the nonlinear coupler. Note that we are using dimensionless units throughout this study.
@@ II.2 Learning and Simulation
The first half of the input time series is used for training and the other half for testing. Note that after measurements at each step, we use the final state of the system before the measurements and after the quantum evolution in the current step to continue with the next step (we do not use the collapsed state after measurements). After training is done, we re-initialize the state of the system to vacuum to start the testing process. For training, Tikhonov regularization is utilized with a regularization parameter that is chosen to get maximum performance. It is worth mentioning that both testing and training were performed in the steady state.
For quantifying the performance of the reservoir, we use the normalized root mean square error (NRMSE), which is calculated as
@@ III Results
Figure 2: Time evolution of entanglement (top row) and corresponding time-series prediction performance (bottom row) at different input frequencies $f$. Top: Behavior of logarithmic negativity in time during testing and training. To ensure the reservoir reflects stable dynamics rather than initial transients, we use the portion of the evolution after the vertical dashed line for training and testing. Bottom: Examples of input sequences at four different frequency scales and The corresponding predictions of the reservoir at coupling rate of $g=0.9$ , input strengths of $\epsilon_{a}=\epsilon_{b}=3$, nonlinearities of $K_{a}=K_{b}=1$ and dissipation rates of $\kappa_{a}=\kappa_{b}=0.1$ in a rotating frame. The performance gets worse as the input frequency increases.
@@ III.1 Input Strength
We randomly selected 30 input strengths ($\epsilon$) from the interval $[0,10]$, while keeping the other parameters constant, including $K_{a}=K_{b}=K=1$ and $g=0.9$, and evaluated NRMSE for input time series at various input frequency scales while computing the time average of the logarithmic negativity in the steady state of the system in each instance. It is important to note that we utilize the same 30 values of input strengths for all frequencies. Furthermore, we ensure that the average excitation number remains below one as discussed in section II.4 and appendix B. The outcome is presented in Fig. 3.
@@ III.2 Kerr Nonlinearity
We randomly selected 30 Kerr nonlinearity values ($K$) from the interval $[0,10]$, while keeping the other parameters constant, including $\epsilon_{a}=\epsilon_{b}=\epsilon=3$ and $g=0.9$, and evaluated the NRMSE for input time series at different input frequency scales while calculating the time average of the logarithmic negativity of the system in each case. Note that we use the same 30 values of $K$ for all frequencies. The results are shown in Fig. 4.
@@ III.3 Coupling Constant
We randomly selected 30 coupling constant values ($g$) from the interval $[0,10]$, while keeping the other parameters constant, including $K_{a}=K_{b}=K=1$ and $\epsilon_{a}=\epsilon_{b}=\epsilon=3$, and evaluated the NRMSE on input time series at different input frequency scales while calculating time average of the logarithmic negativity of the system in each case. Note that we use the same 30 values of $g$ for all frequencies. The results are shown in Fig. 5.
@@ III.4 Entanglement Advantage
Figure 6: Prediction error (NRMSE) vs. entanglement in the case of zero dephasing ($\kappa_{\phi}=0$) and a dissipation rate of $\kappa_{a}=\kappa_{b}=0.1$, for different input frequencies. The data from figs. 3, 4 and 5 are combined and binned based on entanglement levels, quantified via logarithmic negativity, with vertical error bars representing the standard deviation within each bin. Left: Minimum NRMSE within each entanglement bin. No consistent correlation is observed. Middle: Mean NRMSE across all configurations in each bin. This captures the overall performance trend with respect to entanglement. Moderate entanglement improves performance on average. There is an optimum level of entanglement. However, a threshold in the frequency is observed at $f=2.5$ where no trend exists. Right: Maximum NRMSE per entanglement bin. This reflects the worst-performing cases within each bin. A stronger relation is seen here: higher entanglement reduces the worst-case error more than the average error, but again, a frequency threshold is observed. Note: Vertical error bars in all panels represent the standard deviation of the entire data distribution within that entanglement bin. They are included in the Min/Max panels to illustrate the dispersion of the performance data associated with that specific level of entanglement, rather than the uncertainty of the extrema values themselves.
@@ III.5 Role of Dissipation
To examine the impact of dissipation, we repeat the data generation process and binned analysis described earlier in the presence of different dissipation rates, $\kappa_{a}=\kappa_{b}=\kappa$ and compute the average entanglement and NRMSE to produce the same type of plot as the middle panel of Fig. 6, now for different $\kappa$ (see Fig. 9 in Appendix A). It is worth mentioning that our system does not reach a steady state in the absence of dissipation and therefore does not exhibit steady-state entanglement; hence, we do not consider the zero-dissipation case. To better analyze the effect of dissipation, we extract the optimal points (those with minimum error) and plot their corresponding NRMSE as a function of dissipation for different input frequencies, resulting in Fig. 7.
As shown in Fig. 7, a higher dissipation level is better for frequencies below the frequency threshold up to a point where it saturates. By looking at Fig. 9 in Appendix A one can see that realizing an entanglement advantage requires sufficient dissipation, with slightly higher values offering a consistent advantage across all frequency scales below the frequency threshold.
@@ III.6 Role of Dephasing
To examine the impact of dephasing, we repeat the data generation process and binned analysis described earlier in the presence of different dephasing rates $\kappa_{\phi}$, and compute the average entanglement and NRMSE to produce the same type of plot as in Fig. 6, now for different $\kappa_{\phi}$ (Fig. 10 in Appendix A). To more clearly assess the role of dephasing, we identify the optimal points—those corresponding to the minimum prediction error—and plot their associated NRMSE as a function of dephasing for different input frequencies, as shown in Fig.8.
@@
III.7 Nonlinear Prediction Task
To verify the generality of our findings beyond the linear forecasting task, we extended our analysis to a more complex nonlinear prediction task. We applied the 20th-order NARMA kernel to the multi-frequency input signal to generate a highly nonlinear target trajectory. As detailed in Appendix C, our analysis confirms that the key trends observed in the linear task remain robust. Specifically, we find that moderate entanglement continues to be linked to the optimal predictive performance. Furthermore, the frequency threshold effect persists, where the entanglement advantage is lost at higher input frequencies. Interestingly, while dissipation remains beneficial up to a saturation point , we observe that dephasing is strictly detrimental to performance in this nonlinear regime, highlighting the increased sensitivity of nonlinear tasks to phase coherence. Full details and parameter sweeps for this task are provided in Appendix C.
@@ IV Discussion and Conclusion
Our findings indicate that within the various system configurations analyzed, entanglement is indeed linked to improvement in both average and worst-case predictive performance, albeit only within a certain range of operational parameters and up to a threshold in the input frequency. It was noted that optimal average performance, measured in terms of NRMSE, generally takes place when the reservoir maintains a moderate level of entanglement rather than the highest, with the worst-case performance following a similar trend. This effect was most evident for lower input frequency scales, where the temporal structure of the input could be effectively retained and processed by the reservoir dynamics. This kind of relation between quantumness and learning capabilities is consistent with other recent findings in the literature [29, 14, 22], which suggest that excessive “quantumness” may saturate the reservoir or introduce dynamical instability, ultimately hampering learning. However, unlike the spin-network results in [22] which observed an increasing advantage at high frequencies, our coupled-oscillator system exhibits a distinct frequency threshold, which might be related to the excitation number cut-off and the corresponding limit on the average excitation number that we imposed.
Notably, our results demonstrate that the observed connection between entanglement and performance persists under realistic conditions with dissipation and dephasing (please see Appendix A). We showed that increasing dissipation can actually improve performance, up to a point that it saturates. Likewise, we found that the absence of dephasing is generally optimal for performance and entanglement (please see Appendix A); a small amount of dephasing can be tolerated (and at very low input frequency, it even provided a slight benefit).
While our results demonstrate a clear link between entanglement and prediction accuracy, we note that establishing strict causality in physical reservoirs is challenging. However, the robustness of this relation—persisting across different parameters, noise levels, and tasks (including the nonlinear NARMA time-series prediction task, as shown in Appendix C)—suggests that entanglement is linked to the reservoir’s forecasting power in the quantum regime.
There are several limitations to note. First, our study was limited to a two-mode Kerr system with a truncated Hilbert space—a limitation that we addressed by imposing a limit on average excitation number (please see Section II.4, Appendix B). While this setup is well-suited for exploring foundational aspects of QRC, it restricts both the memory depth and the expressive capacity of the reservoir. Second, we focused on a specific class of input signals and employed a standard linear readout with ridge regression. Although this choice facilitates controlled benchmarking, our conclusions may not directly generalize to more diverse or noisy targets or to different output architectures. Additionally, our quantification of entanglement relied on logarithmic negativity, which does not capture all forms of non-classical correlation, such as Wigner negativity, quantum discord, magic/nonstabilizerness, or coherence.
@@
Code Availability
The numerical code used to generate the results in this paper is available on GitHub. Supplementary scripts and datasets can be obtained from the corresponding author upon reasonable request.